one number exceeds another by 6. the sum of the numbers is 66. what are the numbers?
Answer by mananth(16949)

one number exceeds another by 6. the sum of the numbers is 66. what are the numbers?
let thenumber be x
the other number will be x+6
x+6+x=66
2x+6=66
2x=66-6
2x=60
x=30
&36
The numbers are 30 and 36